3. Solve the system of differential equations X= 3x+y y = x+3y
Answer #1

x=3x+y Y = x+3y (2) x=(D-324 (D-3) x = 4 (D-3)² x=(D-3) x = (0-3ly -(1) X (D-3) -(2) (8-6D+9-1) 220 (D-6D+872=0 The auxiliary